OpenStudy (anonymous):

Factor each completely. If the polynomial is not factorable write prime. A. b^2 -12b +36 B. 9c^2 -12c +4 C. p^2 -2p-15 D. c^9 +125

OpenStudy (paxpolaris):

\[(x-y)^2= x^2-2xy+y^2\] this formula works for A and B

OpenStudy (anonymous):

how would I use it ?

OpenStudy (paxpolaris):

A.\[ b^2 -12b +36\\=b^2-2\cdot b \cdot 6+6^2\\=\left( b-6 \right)^2\]

OpenStudy (paxpolaris):

B.\[9c^2-12c+4\\=\left( 3c \right)^2-2\left( 3c \right)\cdot2+2^2\\=\left( 3c-2 \right)^2\]

OpenStudy (anonymous):

could you show me with C. how you do that?

OpenStudy (paxpolaris):

\[p^2-2p-15\\=p^2-5p+3p-15\] you rewrite -2p as -5p +3p .... which still adds up to -2p but multiplies to equal the product of the first and last term: \(-15p^2)

OpenStudy (paxpolaris):

then you factor out what's common in the first two terms and the last two terms\[=p^2-5p+3p-15\\=p \left( p-5 \right)+3\left( p-5 \right)\]

OpenStudy (paxpolaris):

and then you factor out the (p-5)\[\large=\left( p-5 \right)\left( p+3 \right)\]

OpenStudy (anonymous):

and then your left with both of those?

OpenStudy (paxpolaris):

\[p^2 -2p-15\\=(p-5)(p+3)\] yup you have the factored form

OpenStudy (paxpolaris):

for D, use the formula:\[x^3+y^3=(x+y)\left( x^2-xy+y^2 \right)\]

OpenStudy (anonymous):

I really no good with this formula stuff lol sorry what would i do with it after i write it out

OpenStudy (paxpolaris):

\[\large {c^9+125\\=\left( c^3 \right)^3+5^3}\]

OpenStudy (paxpolaris):

\[\large=\left( c^3+5 \right)\left[\left( c^3 \right)^2-\left( c^3 \right)\cdot5+5^2\right]\]

OpenStudy (paxpolaris):

\[\large=\left( c^3+5 \right)\left( c^6-5c^3+25 \right)\] and we can't factor further

OpenStudy (anonymous):

awsome thanks so much :)